OpenStudy (annie96):

What is the equation for a line that passes through the points (-2,0) and (4,12). A) y=-2x+4 B) y=2x+4 C) y=-2x-4 D) y=2x-4 What is the equation for a line that passes through the points (5,-1) and (-10,8). A) y= -3/5x - 2 B) y= 3/5x +2 C) y=-3/5 +2 D) y=3/5x -2

OpenStudy (anonymous):

12/6 =2 so m = 2 y =mx +5 4(12) 12 = 2(4) +b -12 -b B=-4 y=2x-4

OpenStudy (anonymous):

1 is c